### v1

# Rates of Convergence of Generalised Variational Inference Posteriors under Prior Misspecification

Abstract
We prove rates of convergence and robustness to prior misspecification within a Generalised Variational Inference (GVI) framework with bounded divergences. This addresses a significant open challenge for GVI and Federated GVI that employ a different divergence to the Kullback–Leibler under prior misspecification, operate within a subset of possible probability measures, and result in intractable posteriors. Our theoretical contributions cover severe prior misspecification while relying on our ability to restrict the space of possible GVI posterior measures, and infer properties based on this space. In particular, we are able to establish sufficient conditions for existence and uniqueness of GVI posteriors on arbitrary Polish spaces, prove that the GVI posterior measure concentrates on a neighbourhood of loss minimisers, and extend this to rates of convergence regardless of the prior measure.
Rates of Convergence,
keywords:
[class=MSC]
keywords:
[…]
, , and
1 Introduction
The Bayesian, by nature of being Bayesian, can elicit arbitrary priors and may hence encounter prior misspecification; she must either suppose the prior encodes the best available knowledge of the underlying process—this however is scarcely practicable and processes exist for which we cannot intuit priors that lead to consistent posteriors—or accept the arbitrariness [14, 22]. In fact, for any finite number of observations the prior alone may induce any measure a posteriori. Moreover, it is scarcely the case that the model is well specified, that is the model family $\Theta$ will not contain the data generating process [7, 31, 34]. Generalised Variational Inference (GVI) [24] recasts Bayesian posterior updating as an optimisation problem in the calculus of variation over a simpler space of measures, allowing decision makers to mitigate misspecification of both the model and the prior.
The present paper is concerned with characterising the dynamics of updating a posterior measure in the presence of prior and model misspecification through Generalised Variational Inference, as introduced by Knoblauch, Jewson and Damoulas [24]. In particular, we aim to answer the following open problems: under what conditions is a Generalised Variational Inference posterior measure robust to prior misspecification? Do GVI posteriors exist when observations come from an infinite dimensional stochastic process and are these asymptotically consistent? At what rate does this convergence take place?
Asymptotic consistency of posterior measures justifies the updating scheme over its frequentist counterpart by illustrating that we may come arbitrarily close to the optimal given sufficient observations. In Bayesian inference under misspecification asymptotic consistency is often credited when disregarding model and prior misspecification [22, 31, 21]. In fact, Bayesian posteriors concentrate around hypothesis measures in the support of the prior that minimise the Kullback–Leibler divergence rate to the data generating process [30, 7, 5, 15, 32, 33, 22, 16, 31, 34]. This, however, embeds several assumptions which may lead to undesirable posteriors in the misspecified setting: (a) the Kullback–Leibler rate as a choice of loss function, (b) the prior is well specified in that it does not discard hypotheses desired by the decision maker, and leads to consistent posteriors (see e.g. [13, 14]), and moreover, (c) that Bayes is computationally tractable, and hence of use to decision makers.
We already established that (b) may not hold as priors that do not support neighbourhoods around hypotheses with minimal KL divergence rate to the data generating process will lead to concentration around less favourable hypotheses, especially in well specified models, not to mention natural choices of priors that lead to inconsistent posteriors as observed in [13]. Furthermore, since the Kullback–Leibler divergence is known to not be robust (cf. [18]), (a) may lead to suboptimal performance. And as Bayesian inference is rarely computationally tractable, optimising over a simpler space of probability measures—as done in Variational Inference with e.g. the set of Gaussians [19]—allows for computational feasibility. Generalised Variational Inference [24] allows us to mitigate (a), (b), and (c).
[…]
Figure 1: A taxonomy of Generalised Variational Inference. We highlight some special cases.
Suppose $X_{1},X_{2},$... is a sequence of random variables distributed according to some unknown, infinite dimensional data generating process $P_{0}$, and which are modelled through hypothesis measures in the hypothesis space $\Theta$. We begin with some prior measure $\Pi$ and after receiving observations we may update this to $Q_{1},Q_{2},$... accordingly. GVI assumes an optimisation centric viewpoint where we may mitigate (a) by replacing the average negative log–likelihood by an arbitrary loss—such as those based on robust divergences between the data generating process and the hypothesis measure—(b) by replacing the Kullback–Leibler divergence to the prior by an arbitrary divergence, and (c) by optimising over a ‘simpler’ space of measures; see Figure 1 for details. We therefore aim to find a measure within a subset of possible Borel probability measures, $\mathcal{Q}\subset\mathcal{P}(\Theta)$, that achieves
$\inf_{Q\in\mathcal{Q}}\left\{\mathbb{E}_{Q}[n\cdot L(X_{1}^{n},\,\theta)]+\frac{1}{\beta}D(Q:\Pi)\right\}\,.$ (1)
As we are concerned with minimising this functional over the set of possible probability measures, we may write this as
$T_{n}(Q):=n\cdot J_{L_{n}}(Q)+\frac{1}{\beta}D(Q:\Pi)$ (2)
where $J_{L_{n}}(\cdot)$ is the functional given by integration of the loss with $n$ data points against some probability measure. Our results do not require there to be any data, and we may simply drop this dependence whenever convenient. The GVI posterior measure, $Q_{n}$ then minimises this
$Q_{n}\in\operatorname*{arg\,min}_{Q\in\mathcal{Q}}\,T_{n}(Q)$ (3)
whenever it exists. This framework allows decision makers to target a broader set of possible posterior measures, depending on the choices of loss function and divergence, rather than the standard Bayesian posterior; in the case where no misspecification occurs—the prior and likelihood are well specified—the Bayesian posterior is indeed optimal [23, 24, 26, 27]. See Figure 1 for a taxonomy of the GVI framework relating (a), (b), and (c) with the corresponding solutions.
Related Theoretical Work
Rates of convergence of Bayesian posteriors under model misspecification, mild prior misspecification, and highly correlated data are well understood [7, 22, 16, 31]. For generalised Bayesian inference, asymptotic consistency, but not rates of convergence, has been established by Miller [28] in fairly general settings. These concentration results do not apply under severe prior misspecification, requiring the prior to have positive measure on every Kullback–Leibler neighbourhood of the underlying model. Knoblauch [23] established early results on asymptotic consistency of GVI in the parametric setting under fairly strong assumptions, most of which we are able to significantly relax while achieving stronger results.
Main Contributions
This paper develops theory concerning GVI posterior measures over infinite dimensional hypothesis spaces $\Theta$. The main contribution of this paper is Theorem 7 where we establish rates of convergence of GVI posterior measures just slower that $n^{-1}$ when using bounded divergences. In doing so we first present sufficient conditions for existence and uniqueness of GVI posteriors, which we then use as a basis to characterise asymptotic concentration to sets containing the minimisers of the loss. Significantly, our theory allows us to deal with instances in which the prior is severely misspecified, such as in the case where the prior has full measure on a subset of $\Theta$ where the Kullback–Leibler divergence rate is infinite and where the prior and GVI posterior are mutually singular. This is highlighted in Figure 1.
Organisation
In Section 2 we provide the necessary background for Generalised Variational Inference, prior misspecification, and the assumptions used throughout the paper. Section 3 first states and discusses the main results of this paper and we provide the complete proofs of these in Section 5. To highlight our results, we present illustrative examples in Section 4 comparing GVI with Bayesian Inference under prior misspecification in Gaussian location models. In Section 6 we provide some concluding remarks on our results, their relevance to extensions of GVI in Federated Learning [27], and relevant open problems.
2 Preliminaries on Generalised Variational Inference and Prior Misspecification
Let $X_{1},X_{2},...,X_{n}$ be a sequence of observable random variables, short $X_{1}^{n}$, generated by the abstract probability space $(\Omega,\mathcal{F},P_{0})$ and taking values in the measurable space $(\Xi,\mathcal{X})$ with $P_{0}$ distribution. We consider $n\in\mathbb{N}$ and let this tend to infinity for the concentration results. Denote by $\sigma(X_{1}^{n})$ the natural filtration of this sequence. We do not place any restrictions on the data, as long as the loss associated with it satisfies the respective assumptions, especially Assumption 7.
We define our hypothesis measures $P_{\theta}$, with $\theta$ taking values in the measurable space, $(\Theta,\mathcal{T})$ where $\Theta$ is a Polish space, in general infinite dimensional, endowed with its Borel $\sigma$–algebra, and we assume that $P_{0}$ and $P_{\theta}$ admit densities $p_{0}$ and $p_{\theta}$ respectively. As we allow for model misspecification, in the nonparametric setting we suppose that $P_{0}\notin\Theta$ in general [^1] In the parametric setting this translates to $\nexists\theta_{0}\in\Theta$ such that $P_{\theta_{0}}=P_{\theta}$ in the sense of measures.. However, we may find elements of $\Theta$ that are closest, in some sense, to the data generating process $P_{0}$, which we will denote using the ‘’ notation.
Throughout, we will study probability measures in $\mathcal{P}(\Theta),$ the space of all Borel probability measures on $(\Theta,\mathcal{T})$ [20, 11]. We begin with some prior $\Pi$ and update this to the GVI posterior $Q_{n}$.
We further define the terms of Equation 2. Let $D:\mathcal{P}(\Theta)^{2}\rightarrow\mathbb{R}_{\geq 0}$ be a statistical divergence between probability measures. $L(X_{1}^{n},\,\theta)$ is a loss function connecting the data and the hypotheses, and $\mathcal{Q}\subset\mathcal{P}(\Theta)$ is a (sub)space of probability measures we are optimising over. Clearly $n\in\mathbb{N}$ is the number of observables, and the term $\beta>0$ represents a fixed learning rate parameter. We later relax the assumptions on $\beta$.
Notation
We say that eventually $a_{n}\gtrsim b_{n}$ to denote the existence of an unspecified constant $k>0$, depending not on $n$, such that for all $n$ sufficiently large it holds that $a_{n}\geq k\cdot b_{n}$. For a measure $\nu\in\mathcal{P}(\Theta)$ and a set $A\in\mathcal{T}$ we denote $\nu A:=\nu(\{\theta\in\Theta:\theta\in A\})$. Note when using $\subset$, we do not distinguish between strict containment or up to equality as it bears no significance on the theory. We denote $\mathbb{R}_{+}:=\{a\in\mathbb{R}:a>0\}$ and $\mathbb{R}_{\geq 0}:=\{a\in\mathbb{R}:a\geq 0\}$.
2.1 Generalised Variational Inference and Related Methods
The Bayesian posterior—consisting of the likelihood function of the data, $p_{\theta}(X_{1}^{n})$, and a prior measure, $\Pi$—is
$\Pi_{n}A:=\cfrac{\int_{A}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}{\int_{\Theta}p_{\theta}(X_{1}^{n})\,\Pi(d\theta)}$
as usual. If we replace the likelihood by the exponential of a negative loss acting on the data we get Generalised Bayesian Inference (GBI) [9, 28], or a Gibbs posterior [2]. If we discard the prior, we may instead get an energy–based, frequentist model. As shown in [36] Bayes may be equivalently formulated as an optimisation problem where $\Pi_{n}$ is the posterior minimising the ‘Evidence Lower Bound’—the expectation of the negative log–likelihood, and the Kullback–Leibler divergence to the prior—over all probability measures acting on the hypothesis space. GVI instead replaces the negative log–likelihood by any loss function, similar to GBI, as well as the KL divergence by any divergence, generalising these previous approaches. Tangentially, minimising a divergence other than the KL divergence to the Bayesian posterior, not just the prior, may result in nicer optimisation problems in practice, e.g. in Rényi variational inference [25]. This however does not nicely decompose into an expectation acting on the loss and a divergence measure acting on the prior, conflating model and prior misspecification; see also section 2.3.2 in [24]. Typical choices of loss functions are those that either offer attractive computational benefits, and those that incorporate the hypothesis measure robustly, including those based on robust divergences [27].
2.2 Prior Misspecification
As identified in [34, 31] the elements of $\Theta$ that the Bayesian posterior targets are those minimising the Kullback–Leibler divergence between the data generating process and the hypothesis measure. Denoting this as a function on $\Theta$, we write:
$h(\theta)=\lim_{n}\frac{1}{n}\mathbb{E}_{P_{0}}\left[\log\frac{p_{0}(x_{1}^{n})}{p_{\theta}(x_{1}^{n})}\right]$
which may possibly be infinite; note that the expectation is with respect to $P_{0}$ and hence over $\Omega$. This is the loss function targeted in classical Bayesian inference, in particular the elements of $\Theta$ that minimise the above minimise the Kullback–Leibler divergence between the data generating process and the hypothesis measure. If the prior has full measure on the set of elements where this rate is infinite, denote this as $I:=\{\theta\in\Theta:h(\theta)=\infty\}$, then all previous results on consistency, let alone rates of convergence, of (Generalised) Bayesian inference fail [31]. In this case we say that the prior is severely misspecified. In this we also include priors that lead to inconsistent posteriors such as those observed in [13].
To make this rigorous, we place the following definition of robustness to prior misspecification.
Definition 1.
Let $\mathcal{G}\subset\mathcal{P}(\Theta)$ be the set of all possible priors. We say a posterior is asymptotically robust to prior misspecification w.r.t. $\mathcal{G}$ if the posterior is consistent for all priors $\Pi\in\mathcal{G}$.
Classical results on Bayesian sensitivity analysis deal with small perturbations of some reference prior by some other measure through additive or multiplicative misspecification [12, 17]. That is some reference prior $\Pi$ is perturbed by another probability measure $G\in\mathcal{P}(\Theta)$ through a constant $\varepsilon\in[0,1]$, e.g. $\Pi_{\varepsilon}=(1-\varepsilon)\Pi+\varepsilon G$. Our results are different to this line of work, and we will not use such explicit formulations for $\mathcal{G}$. We discuss this distinction in Appendix B.
2.3 Illustrative Example: Bayesian Inference under Prior Misspecification
To illustrate the non–robustness of Bayesian inference under prior misspecification, we give a simple, parametric example with $\Theta=\mathbb{R}$. Consider the classical Bayesian posterior, and assume that $p_{0}$ is Gaussian with unknown mean and known variance, $X_{i}\sim\mathcal{N}(\theta_{0},\sigma_{p}^{2})$. The prior is a contaminated Gaussian, $\pi\sim\mathcal{N}(\mu_{\pi},\sigma_{\pi}^{2})$ where $\mu_{\pi}\in\mathbb{R}$. Then the posterior on the mean, $\theta$, is Gaussian and characterised by the following variance and mean:
$\sigma_{n}^{2}=\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)^{-1},\,$ $\mu_{n}=\sigma_{n}^{2}\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}\frac{X_{i}}{\sigma_{p}^{2}}\right)$
For fixed $n$, and given knowledge of the data, we may pick $\mu_{\pi}$ and $\sigma_{\pi}$ to achieve any Gaussian distribution a posteriori, disregarding the data entirely.
Moreover, the Kullback–Leibler divergence to the prior has closed form:
$D_{KL}(q:\pi)=\log\frac{\sigma_{\pi}}{\sigma_{q}}+\frac{(\mu_{q}-\mu_{\pi})^{2}+\sigma_{q}^{2}}{2\sigma_{\pi}^{2}}-\frac{1}{2}$
which is convex in both $\mu_{q}$ and $\sigma_{q}$. Therefore, as $\mu_{\pi}\rightarrow\infty$, this increases, and hence forces $\mu_{q}$ to match the prior mean as the gradient is minimised if and only if their means are equal. See Appendix C for further details.
2.4 Assumptions
For the GVI objective to make sense we require some standard assumptions that will ensure the problem is not pathological.
Assumption 1.
The losses $L(X_{1}^{n},\,\theta)$ are $\sigma(X_{1}^{n})\times\mathcal{T}$–measurable and lower bounded [^2] Loss functions used in ML are typically non–negative, we, in accordance with the literature on GVI and GBI, allow for more general functions through the lower boundedness. for any $n\in\mathbb{N}$.
This standard assumption ensures that the GVI objective is not pathological, making sense as a machine learning objective. Allowing the loss to be $\infty$’ would imply that we could pick a Dirac–delta measure at any such point. However, we do not place any restrictions on the underlying data distribution, and strictly speaking, our results do not require there to be any data at all; we could simply work with a sequence of $\mathcal{T}$–measurable functions as in [28]. Where this distinction is not necessary, to lighten notation, we may write $L_{n}(\theta):=L(X_{1}^{n},\,\theta)$.
Furthermore, we require that the subspace of “simpler” measures we are optimising over is well behaved and in particular that we are able to find an element of this space for which the GVI functional is finite.
Assumption 2.
There exists at least one $Q\in\mathcal{Q}$ such that $T_{n}(Q)<\infty$.
Further assumptions on this space $\mathcal{Q}$ such as it being closed in the weak topology or being convex are made explicit in the theorems where required. These are sufficient to ensure existence and uniqueness of GVI posteriors as soon as there exists a single element of $\mathcal{Q}$ for which Equation 2 is finite (see Theorems 2, 4 and 8).
As we desire a framework that is able to deal with severe prior misspecification, we require the following assumption on the divergence.
Assumption 3.
For given $\mathcal{Q}$ and $\mathcal{G}$ the divergence in Equation 2 is bounded above, i.e. $\exists M>0$ such that $\forall Q\in\mathcal{Q}$ and $\forall\Pi\in\mathcal{G}$ we have $D(Q:\Pi)\leq M$.
The assumption here is a stricter requirement and violated by many common divergences, such as the KL divergence, as shown in the previous section. However such bounded–above divergences do exist, a prime example being the Total Variation Distance. This assumption allows us to bound the influence of the prior on the GVI objective, and forces the GVI posterior to not discard the data. We find that this allows for favourable behaviour of GVI in the presence of prior misspecification, ensuring frequentist consistency under very general assumptions, including the case where the GVI posterior and the prior are mutually singular.
Remark.
It is worth noting that this may be satisfied by both bounded divergences, such as the total variation distance, as well as specific choices of the set of measures we are optimising over, $\mathcal{Q}$, and the set of possible priors, $\mathcal{G}$. In the former case, we may allow for $\mathcal{G}=\mathcal{P}(\Theta)$ and $Q\subset\mathcal{P}(\Theta)$ without restriction. The latter is more restrictive and includes the case where $\mathcal{G}$ is a singleton and $Q$ the sublevel sets of $\mathcal{P}(\Theta)$.
The next assumption is of a technical nature, saying that we may achieve the infimum of the loss functional over the set $\mathcal{Q}$ with some element from that set.
Assumption 4.
$\exists P_{n}^{\star}\in\mathcal{Q}$ such that $\forall Q\in\mathcal{Q}$ it holds that $J_{L_{n}}(P_{n}^{\star})\leq J_{L_{n}}(Q)$.
This is not a particularly restrictive assumption, and will generally be satisfied by typical choices of $\mathcal{Q}$. We use the ‘’ notation throughout to denote minimisers, or a relation to a minimiser.
To prove existence and uniqueness of the GVI posterior, we require additional assumptions on the loss and the divergence.
Assumption 5.
The losses as functions on $\Theta$, $L_{n}:\Theta\rightarrow\mathbb{R}$, are coercive and lower semi–continuous for all $n\in\mathbb{N}$.
This assumption is crucial to establish the existence of minimisers as we require coerciveness of the loss in order to establish tightness of a sequence of probability measures. The lower semi–continuity of the loss is easier to verify and the same then immediately follows for the functional [11, 4], however it is slightly stronger than required as we merely need the lower semi–continuity of the functional $J_{L_{n}}(\cdot)$.
Remark.
This assumption carries several implications. The lower semi–continuity of the loss allows it to achieve its infima. If $\mathcal{Q}$ is weak closed then we may discard Assumption 4 as such a measure will necessarily exist. Note also that the coercivity of the loss may be discarded if $\Theta$ were compact.
Assumption 6.
The divergence $D:\mathcal{P}(\Theta)\times\mathcal{P}(\Theta)\rightarrow\mathbb{R}_{\geq 0}$ is lower semi–continuous in its first argument.
In combination with Assumption 5, this establishes that the entire GVI functional is lower semi–continuous, a key tool in proving the existence of minimisers to Equation 3, and is satisfied by key divergences, such as the KL divergence or the total variation distance.
Finally, in order to reason about the asymptotic properties and rates of convergence of the GVI posteriors we need that:
Assumption 7.
There exists a $\mathcal{T}$-measurable function $L:\Theta\rightarrow\mathbb{R}$ such that $\forall\theta\in\Theta$
$\lim_{n\rightarrow\infty}L(X_{1}^{n},\,\theta)=L(\theta)$
$P_{0}$ almost–surely, which is lower semi–continuous and that it $P_{0}$–a.s. holds that
$\inf_{\theta\in\Theta}L(\theta)\geq\limsup_{n}\inf_{\theta\in\Theta}L(X_{1}^{n},\theta)>-\infty\,.$
This assumption is required for the concentration results of Theorems 6 and 7, and is standard in Bayesian asymptotics (see e.g. [31]) where an asymptotic equipartition property, as in the Shannon–MacMillan–Breimann theorem, is required to hold. Unless otherwise stated all convergence takes place in the number of observables $n$.
As we will be working with the infima of the different loss functions throughout, we follow [31] by adopting further notational abbreviations stemming from large deviation theory.
$L_{n}(A)\equiv\inf_{\theta\in A}\,L(X_{1}^{n},\,\theta)\,,\qquad L(A)\equiv\inf_{\theta\in A}\,L(\theta)$ (4)
As we have not fixed a metric, we follow [7, 31, 28] and use the following set as a measure of closeness:
$N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}\,.$ (5)
3 Main Results
We first provide the statements of the main theoretical results. We begin by characterising the GVI posteriors under bounded divergences through the empirical loss minimiser, find sufficient conditions for existence and uniqueness in infinite dimensional hypothesis spaces, and prove asymptotic consistency, culminating in a rates of convergence result. Finally, we briefly touch upon GVI posteriors under unbounded divergences, dispensing of Assumption 3. The proofs are provided in the corresponding Sections 5.1, 5.2, 5.3, 5.4, 5.5 and 5.6.
3.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
Any prior within the set $\mathcal{G}\subset\mathcal{P}(\Theta)$ cannot have infinite influence under Assumptions 3 and 4, and therefore GVI prefers measures that are close to the empirical loss minimiser, such that the prior has limited influence if it does not agree with the data.
Theorem 1.
Under Assumptions 1, 2, 3 and 4, for all priors $\Pi\in\mathcal{G}$, the respective GVI posterior $Q_{n}$ of Equation 3 is in the set
$\mathcal{R}_{n}^{\star}:=\left\{Q\in\mathcal{Q}:\int_{\Theta}L(X_{1}^{n},\,\theta)\left[Q(d\theta)-P_{n}^{\star}(d\theta)\right]\leq\frac{M}{n\beta}\right\}.$
Moreover, suppose that $\mathcal{Q}$ is closed with respect to weak convergence and that the loss satisfies Assumption 5, then $\mathcal{R}^{\star}_{n}$ is compact with respect to weak convergence.
Significantly, this result is independent of the choice of prior from the set $\mathcal{G}$. It implies that the GVI posterior may not deviate too far from the empirical loss minimising measure, $P_{n}^{\star}$, albeit in a very rough ‘neighbourhood’. This result implies that the divergence term may never skew posteriors away from the data, given such bounded divergences, even under severe prior misspecification where the traditional Kullback–Leibler divergence were infinite. It is not even necessary for the support of the prior to intersect a subset of $\Theta$ on which the GVI posterior has full measure, nor the family $\mathcal{Q}$ and the prior $\Pi$ to be dominated by some common reference measure.
The weak compactness result allows us to characterise the existence and uniqueness of GVI posteriors. If $\Theta$ were compact then the weak compactness of $\mathcal{R}_{n}^{\star}$ above would be immediate, as any set of measures on $\Theta$ will be tight. However, as $\Theta$ will rarely be compact, we further require that the loss is coercive in order to achieve the same result.
3.2 Existence and Uniqueness of a Generalised Variational Inference Posterior
Having established that $\mathcal{R}_{n}^{\star}$ is compact with respect to weak convergence of probability measures, we are able to demonstrate the existence of minimisers of the GVI functional, Equation 3. We prove existence and uniqueness of GVI posteriors over arbitrary Polish spaces as soon as there is a single element in $\mathcal{Q}$ for which Equation 3 is finite, extending the previous results of Knoblauch [23] and Wild et al. [35]. In particular, we allow for $\Theta$ to be an infinite dimensional Polish space, rather than a finite dimensional, complete, and separable normed space. Far less do we require for $\mathcal{Q}$ to be convex; this is only required to establish uniqueness.
Theorem 2.
Make Assumptions 1, 2, 3, 4, 5 and 6 and suppose also that $\mathcal{Q}$ is closed with respect to weak convergence. Then the GVI minimiser $Q_{n}$, of Equation 3, exists.
Establishing the existence of GVI posterior measures provides that the GVI objective is in fact well specified and a reasonable framework for decision makers in the nonparametric setting. While the above theorem relies on the fact that $\mathcal{R}_{n}^{\star}$ is weak compact, and hence on the boundedness of the divergence, we later establish that this assumption is not necessary given that the prior is fixed. We provide this result in Section 3.6 in the form of Proposition 8.
Unfortunately, as the set $\mathcal{Q}$ is not convex in typical situations of interest, we generally do not have uniqueness of the GVI posterior (see [29] for non–uniqueness under Kullback–Leibler approximations). For instance, the set of Gaussian measures is not convex, nor are typical mixture models [29]. However, we can make the following observation:
Corollary 3.
A convex combination of elements from $\mathcal{R}^{\star}_{n}$ is in $\mathcal{R}^{\star}_{n}$ if and only if that convex combination is also in $\mathcal{Q}$. That is, pick $\mu,\nu\in\mathcal{R}_{n}^{\star}$ and some $a\in[0,1]$, then
$\gamma:=a\mu+(1-a)\nu\in\mathcal{R}_{n}^{\star}\iff\gamma\in\mathcal{Q}\,.$
It follows as special cases, that (a) if $\mathcal{Q}$ is convex then so is $\mathcal{R}_{n}^{\star}$, and (b) if $\mathcal{Q}$ is a set of mixtures, not necessarily convex, then if individual components are contained within $\mathcal{R}_{n}^{\star}$, then so are all permitted mixtures between these. If $\mathcal{Q}$ is indeed convex then the following, stronger result on the uniqueness of the GVI minimiser then immediately follows:
Corollary 4.
Let the assumptions be as in Theorem 2. Suppose further that the divergence is strictly convex in its first argument and that $\mathcal{Q}$ is convex. Then there exists a unique minimiser of the functional $\nu\mapsto T_{n}(\nu)$.
Therefore, under the conditions of the corollary above, the GVI minimiser will be an extreme point of not only $\mathcal{R}_{n}^{\star}$ but also of $\mathcal{Q}$, exists, and is unique.
3.3 Support of the Prior is not required
Significantly, our results are independent of the choice of prior $\Pi\in\mathcal{G}$. This implies that the support of the prior need not contain the support of the GVI posterior, hence assumptions such as those in [28] where we require the prior to place positive weight on a neighbourhood of the expected loss minimiser are not necessary. Therefore, even if the prior is severely misspecified is its influence may not skew the posterior too far away from the data.
Proposition 5.
Theorem 1 holds even if $\operatorname*{supp}(Q_{n})\cap\operatorname*{supp}(\Pi)=\emptyset$.
That is, the prior and the GVI posterior may be mutually singular. Therefore, we get a stronger result than in classical and generalised Bayesian/Variational Inference that require for the prior to have positive mass on the support of the posterior measure. Since Assumption 3 applies, this will typically only be the case when the divergence is a bounded divergence rather then when $\mathcal{Q}$ and $\mathcal{G}$ are restricted. Nevertheless, this allows us to overcome challenges, as discussed in [13, 14], where Bayesian posteriors are inconsistent. Hence leading to robustness to severe prior misspecification.
This proposition is significant as it will imply that the prior need not put positive measure on Kullback–Leibler neighbourhoods of the data generating process, $P_{0}$. It therefore allows for the consistency and rate of convergence results that follow to hold in severely prior misspecified setting, even when the prior has no measure on open sets of the hypothesis space close to the loss minimisers, in the sense of $N_{\varepsilon}$.
3.4 Asymptotics and Frequentist Consistency
Theorem 1 allows us to study the asymptotics of the GVI posterior without requiring us to explicitly know it. By Theorem 2 we additionally know that the sequence of GVI posteriors exist, so we may study their asymptotic behaviour. In key difference to previous work on (generalised) Bayesian posteriors [28, 22, 31] we do not require explicit knowledge of the GVI posterior. Furthermore, when comparing our results to the frequentist consistency result of [23], the assumptions on the loss (and the data) are far less strict, and we do not require $\Theta$ to be a finite dimensional separable normed space. While the bounded divergence is a stricter requirement this additionally allows for robustness to prior misspecification as in Definition 1. To reason about the asymptotics of the sequence of GVI posterior measures we need that these exist, so we suppose in the remaining that $\mathcal{Q}$ is closed with respect to weak convergence.
A key challenge here is that we require for the family $\mathcal{Q}$ to be rich enough, that is the measures that we desire to concentrate on exists within $\mathcal{Q}$ a priori. Naturally, this challenge is non–existent in consistency results of Bayesian posteriors as $\mathcal{Q}=\mathcal{P}(\Theta)$, we however need to be able to deal with subsets of $\mathcal{P}(\Theta)$. We solve this issue by requiring for $\mathcal{Q}$ to allow for Dirac measures; by this we mean that $\mathcal{Q}$ is closed w.r.t. weak convergence of probability measures, and that for all $n$ sufficiently large it holds $P_{0}$–a.s. that $\inf_{\nu\in\mathcal{Q}}\,J_{L_{n}}(\nu)=L_{n}(\Theta)$. Note that we defined $L_{n}(A)\equiv\inf_{\theta\in A}L_{n}(\theta)$ in Equation 4.
Theorem 6.
Make Assumptions 1, 2, 3, 4, 5, 6 and 7. Suppose further that $\mathcal{Q}$ allows for Dirac delta measures, and that $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ we almost surely have that
$\liminf_{n}L_{n}(A)>L(\Theta)\,.$
Then $P_{0}$–a.s. for any such $A\in\mathcal{T}$
$Q_{n}A\rightarrow 0.$
Moreover, suppose there exists a unique $\theta^{\star}\in\Theta$ minimising $L(\theta)$, that is $L(\theta^{\star})=L(\Theta)$ and that $\forall\theta^{\prime}\in\Theta$ where $\theta^{\prime}\neq\theta^{\star}$ $L(\theta^{\prime})>L(\theta^{\star})$, then $P_{0}$–a.s.
$Q_{n}\overset{{w}^{\star}}{\longrightarrow}\delta_{\theta^{\star}}\,.$
This result implies concentration of the GVI posterior on a subset of $\Theta$ on which the converged loss $L$ is minimised. In fact we may pick $A$ such that we concentrate on its complement, in particular $A^{c}=N_{\varepsilon}:=\{\theta:L(\theta)\leq L(\Theta)+\varepsilon\}$ for any $\varepsilon>0$.
In the moreover part we show the weak ($w^{\star}$) convergence of $(Q_{n})$ to $\delta_{\theta^{\star}}$, which we may interpret as the expected loss minimiser of $L(\theta)$, i.e. $P^{\star}_{\infty}$. The extension to infinite dimensional Polish spaces and the weakening of assumptions make this is a considerable improvement over the previous result on weak convergence in [23].
3.5 Rate of Convergence
As our main result of the paper, we are able to derive a rate of convergence result in the number of observations $n$. In fact, we may pick a non–increasing sequence $\varepsilon_{n}$ decaying slowly enough such that the concentration still holds on sets $N_{\varepsilon_{n}}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon_{n}\}\,.$
Theorem 7.
Make Assumptions 1, 2, 3, 4, 5, 6 and 7, and suppose $\mathcal{Q}$ allows for Dirac delta measures. Pick a positive sequence $\varepsilon_{n}$ such that $\varepsilon_{n}\rightarrow 0$ and $n\varepsilon_{n}\rightarrow\infty$. If
$L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta)\gtrsim\varepsilon_{n}$
eventually almost surely, then
$Q_{n}N_{\varepsilon_{n}}\rightarrow 1$
$P_{0}$–a.s.
Significantly, this is the first result that establishes rates of convergence for Generalised Variational Inference posterior measures. Here we allow for sequences $\varepsilon_{n}$ that converge just slower than $n^{-1}$ which should be compared to the result of [31] who finds the same rates for Bayesian posteriors with highly correlated data using Kullback–Leibler neighbourhoods. In [16] the authors finds rates of order $n^{-1/2}$ for the Bayesian posterior using stronger definitions of $N_{\varepsilon_{n}}$ through the Hellinger metric.
Moreover, since this result builds on Theorem 1 it follows that Proposition 5 applies and that the above theorem is independent of the choice of prior $\Pi\in\mathcal{G}$. This implies that we do not require $\Pi(N_{\varepsilon_{n}})>0$ for any $n\in\mathbb{N}$, and the GVI posterior may hence concentrate on hypotheses that minimise the converged loss, $L$, rather than being restricted to the support of the prior which may be misspecified.
3.6 Extensions
Throughout, we have assumed that Assumption 3 holds, that is the divergence between all possible posteriors, $\mathcal{Q}$, and all priors, $\mathcal{G}$, is bounded. This has allowed us to make significant progress in establishing asymptotic consistency and rates of convergence of GVI posterior measures; under unbounded divergences this is an open problem.
Remark.
Assumption 3 can also be satisfied by restricting the set of possible posterior measures, $\mathcal{Q}$, to the sublevel sets of an unbounded divergence to some prior:
$\mathcal{Q}|_{M}:=\{\nu\in\mathcal{Q}:D(\nu:\Pi)\leq M\}.$
Theorem 1 may now be satisfied with $\mathcal{Q}|_{M}$ and $\mathcal{G}=\{\Pi\}$. This, however, does not guarantee robustness to prior misspecification in the same generality as using a bounded divergence, due to the set $\mathcal{G}$ from Definition 1 being smaller. In fact, $\mathcal{Q}|_{M}$ implicitly makes assumptions on the absolute continuity of its members to the prior.
Here, we explore to what extend we may extend our results to unbounded divergences. In particular, we find that we may establish existence and uniqueness without Assumption 3 through additional work, and we make some progress on asymptotic consistency by replacing Assumption 3 by different assumptions on the divergence. Nevertheless, rates of convergence of GVI posterior measures under unbounded divergences remain an open problem.
3.6.1 Existence and Uniqueness under Unbounded Divergences
As long as there is a single element of $\mathcal{Q}$ for which the GVI functional is finite, then the posterior exists. If this further holds for all priors within the subspace $\mathcal{G}$, then it exists regardless of the prior measure.
Proposition 8.
Make Assumptions 2, 1, 6 and 5 and suppose that $\mathcal{Q}$ is closed with respect to weak convergence. Then the GVI posterior minimising Equation 3 exists.
Moreover, suppose also that the divergence is strictly convex in its first argument and that $\mathcal{Q}$ is convex. Then the GVI posterior exists and is unique.
We have now established sufficient conditions for existence and uniqueness of Generalised Variational Inference posteriors under unbounded divergences over arbitrary Polish spaces. Significantly, a GVI posterior will exist as soon as a single measure within $\mathcal{Q}$ has finite GVI objective.
3.6.2 Relaxation of $M$ and $\beta$
There is no objective reason why the different inputs into the GVI objective should be fixed with respect to incoming observations. In fact, Theorems 6 and 7 remain valid even if $\beta$, $M$, $\Pi$, or $D$ vary with $n$. We simply require that the respective assumptions are satisfied for each $n$, but these may affect the rate of convergence.
Proposition 9.
Let $\beta^{-1}:\mathbb{N}\rightarrow\mathbb{R}_{+}$ and $M:\mathbb{N}\rightarrow\mathbb{R}_{+}$ be functions on $n$ mapping to strictly positive numbers. If
$M(n)\beta^{-1}(n)n^{-1}\rightarrow 0,\quad n\rightarrow\infty$
then the conclusion of Theorem 6 is unchanged. Further, Theorem 7 holds with rates
$\varepsilon_{n}>0,\qquad\varepsilon_{n}\rightarrow 0,\qquad n\varepsilon_{n}M(n)^{-1}\beta(n)\rightarrow\infty.$
Remark.
This for instance allows for the following divergence:
$D_{TV}^{(\sqrt{n})}(Q:\Pi):=n^{1/2}||Q-\Pi||_{TV},$
preventing over–concentration around the data in cases where the prior is not severely misspecified while maintaining the results on rates of convergence. We may hence achieve an interplay between Bayesian and frequentist inference.
3.6.3 Concentration of GVI Posteriors under Arbitrary Divergences
For completeness, we also include a result on asymptotic consistency of infinite dimensional GVI posteriors under unbounded divergences. Unfortunately, this requires restrictive assumptions on the set $\mathcal{Q}$, and requires that the loss and the divergence are well behaved. Specifically, we require that a sequence of measures exists which concentrate around loss minimisers, and that eventually the contribution of the divergence of these measures is of order $\mathcal{O}(n^{\alpha})$ for $\alpha<1$. Nevertheless, if this is satisfied we can conclude that the sequence of GVI posteriors concentrates on sets containing loss minimisers.
Theorem 10.
Make Assumptions 1, 2, 5, 6 and 7. Suppose also that there exists a sequence $(\nu_{n})\subset\mathcal{Q}$, not necessarily GVI minimisers, for which (i) $(J_{L_{n}}(\nu_{n})-L_{n}(\Theta))\rightarrow 0$ $P_{0}$–a.s., (ii) $n^{-1}D(\nu_{n}:\Pi)\rightarrow 0$, and (iii) $\forall A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ it almost surely holds that $\liminf_{n}L_{n}(A)>L(\Theta)$. Then, for all such sets $A$,
$Q_{n}A\rightarrow 0$
with $P_{0}$–probability 1.
Here, in key difference to Theorem 6, we require an explicit choice of prior $\Pi$, and while this result allows us to reason about the concentration of arbitrary GVI posteriors on infinite dimensional Polish spaces $\Theta$, it unfortunately tells us nothing about the rate of convergence.
4 Illustrative Examples
We provide worked examples in support of the theory.
4.1 Existence of Everywhere Bounded Divergences
The prime example of such divergences, that satisfy Assumption 3 for unrestricted choices of $\mathcal{Q}$ and $\mathcal{G}$, is the Total Variation Distance defined as:
$D_{TV}(Q:\Pi):=\sup_{A\in\mathcal{T}}\left|QA-\Pi A\right|\leq 1.$ (6)
Since $Q$ and $\Pi$ are probability measures, which are by definition non–negative, the upper bound is achieved for instance in the case when $Q$ and $\Pi$ are mutually singular measures.
As an immediate consequence, the Hellinger Distance is also bounded above since by a well known result, see for instance [29, Lemma B.5],
$D_{H}^{2}(Q:\Pi)=\int_{\Theta}\left(\sqrt{\frac{dQ}{d\mu}(\theta)}-\sqrt{\dfrac{d\Pi}{d\mu}(\theta)}\right)^{2}d\mu(d\theta)\leq D_{TV}(Q:\Pi).$ (7)
Furthermore, such a divergence can be constructed as part of the class of ff–divergences [1, 3]. We follow [3] and define these as
$D_{f}(Q:\Pi)=\mathbb{E}_{Q}\left[f\left(\frac{d\Pi}{dQ}\right)\right]$
for some convex function $f\in C^{1}(\mathbb{R}_{\geq 0},\mathbb{R})$, such that $f(1)=0$. This is a vast class of divergences and contains for instance the KL and the $\chi^{2}$ divergences.
Making use of a result in [10], we know that for ff–divergences, there is an upper bound whenever the following limit exists and is finite:
$0\leq D_{f}(Q:\Pi)\leq\lim_{u\rightarrow 0^{+}}\left\{f(u)+uf\left(\frac{1}{u}\right)\right\}$ (8)
Note that this is independent of any choice of Probability measures $Q$ and $\Pi$, and solely reliant on the function $f$.
For instance, we can pick the Le Cam ff–divergence, which uses
$f(u)=\frac{(u-1)^{2}}{u+1}$
for which this limit exists and, through basic analysis, is 2.
4.2 Gaussian Location Model
Continuing from Section 2.3, we use the example of a Gaussian likelihood, and some Gaussian prior with fixed unit variance but misspecified mean. For simplicity, we assume that $X_{1}^{n}\overset{iid}{\sim}P_{0}:=\mathcal{N}(\theta_{0},1)$ and that the variance is known. In terms of densities, to estimate this normal location model, we place a Gaussian prior $\pi(\theta)\sim\mathcal{N}(\mu_{\pi},1)$ over the possible location parameters in $\Theta=\mathbb{R}$, and we assume that the likelihood hypothesis is $p_{\theta}(X)\sim\mathcal{N}(\theta,1)$. The standard loss for Bayesian inference is then naturally
$L(X_{1}^{n},\,\theta)=-\frac{1}{n}\log p_{\theta}(X_{1}^{n})=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(X_{i})$
which $P_{0}$–a.s. converges in $n$ to its limit, $L(\theta)=\mathbb{E}_{X\sim P_{0}}[-\log p_{\theta}(X)]$. The distribution $P^{\star}_{n}$ in $\mathcal{Q}$, the set of Gaussian measures, is then given by:
$p^{\star}_{n}(\theta)=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\mathbb{E}_{q(\theta)}\left[-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(X_{i})\right]$
which we can evaluate for some $q\sim\mathcal{N}(\mu,\sigma^{2})$ as
$-\frac{1}{n}{{\int}}q(\theta)\left(\sum_{i=1}^{n}\log\left[\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{(X_{i}-\theta)^{2}}{2}\right\}\right]\right)\,d\theta$
$=\underbrace{{{\int}}q(\theta)\frac{1}{2}\log 2\pi\,d\theta}_{=\frac{1}{2}\log 2\pi}+\frac{1}{n}\int q(\theta)\underbrace{\left(\sum_{i=1}^{n}\frac{(X_{i}-\theta)^{2}}{2}\right)}_{=\frac{{\sum_{i}X_{i}^{2}-2\theta\sum_{i}X_{i}+n\theta^{2}}}{2}}\,d\theta$
$=\frac{1}{2}\log 2\pi+\sum_{i=1}^{n}\frac{X_{i}^{2}}{2n}-\frac{\mu\sum_{i=1}^{n}X_{i}}{n}+\frac{(\mu^{2}+\sigma^{2})}{2}$
which is convex in both $\mu$ and $\sigma$, and the minimiser corresponds to the degenerate Gaussian centred at the empirical mean, $\overline{X_{1}^{n}}$. So we let $P^{\star}_{n}(\theta)=\delta_{\overline{X_{1}^{n}}}(\theta)$, and find that
$J_{L_{n}}(P_{n}^{\star})=\frac{1}{2}\log 2\pi+\frac{1}{2n}\sum_{i=1}^{n}X_{i}^{2}-\frac{1}{2}\left(\overline{X_{1}^{n}}\right)^{2}.$
We can then find the set $\mathcal{R}^{\star}_{n}$ in which the GVI posterior may lie.
$\int_{\Theta}L(X_{1}^{n},\,\theta)\left[Q(d\theta)-P_{n}^{\star}(d\theta)\right]\leq\frac{M}{\beta n}$
$\implies\left({\frac{1}{2}\log 2\pi}+\sum_{i=1}^{n}\frac{X_{i}^{2}}{2n}-\frac{\mu\sum_{i=1}^{n}X_{i}}{n}+\frac{(\mu^{2}+\sigma^{2})}{2}\right)-J_{L_{n}}(P_{n}^{\star})\leq\frac{M}{\beta n}$
$\implies\left(\frac{1}{2}\overline{X_{1}^{n}}-\mu\right)\overline{X_{1}^{n}}+\frac{\mu^{2}+\sigma^{2}}{2}\leq\frac{M}{\beta n}$ (9)
We note that the LHS is strictly non-negative, and by elementary analysis this parabola has at least one real root, therefore:
$\mu^{2}-2\mu\overline{X_{1}^{n}}-\frac{2M}{\beta n}+\left(\sigma^{2}+\overline{X_{1}^{n}}^{2}\right)\leq 0$
$\implies$ $\mu\in\overline{\mathrm{B}\left(\overline{X_{1}^{n}},\;\left(\frac{2M}{\beta n}-\sigma^{2}\right)\right)}$ (10)
$\implies$ $\sigma^{2}\in\left[0,\;\mu^{2}-2\mu\overline{X_{1}^{n}}+\overline{X_{1}^{n}}^{2}-\frac{2M}{\beta n}\right]$ (11)
where $\overline{\mathrm{B}(x,r)}$ denoted the closed ball of radius $r$ centred at $x$. Hence, the space of possible GVI posteriors is compact regardless of the prior mean $\mu_{\pi}$, in difference to the Bayesian posterior found in Section 2.3.
Furthermore, since the LHS in Equation 9 is non–negative, it is straightforward to see that we converge to the mean of the DGP as $n\rightarrow\infty$, i.e. $\lim_{n}\frac{1}{n}\sum_{i}^{n}X_{i}=\mathbb{E}_{P_{0}}[X]=\theta_{0}$, with parameters $\mu_{\infty}=\theta_{0}$ and $\sigma_{\infty}^{2}=0$ $P_{0}$ almost surely.
This happens irrespective of the prior, even in the case where $\pi(\theta)=\delta_{\theta_{\pi}}$ and $\theta_{\pi}\rightarrow\infty$, or when $\theta_{0}\notin\operatorname*{supp}(\Pi)$. This result stands in contrast to concentration results of GBI posteriors such as Theorem 2 of [28].
4.2.1 A Comparison with Bayesian Inference
To highlight this relevance, we compare the set $\mathcal{R}^{\star}_{n}$ with the Bayesian posterior under prior misspecification and examine in which set the prior may lie in order to produce the worst case posterior found in $\mathcal{R}_{n}^{\star}$.
From Section 2.3, we know that the Bayesian posterior mean and variance are given as
$\sigma_{B}^{2}=\left(\sigma_{\pi}^{-2}+n\right)^{-1},\,$ $\mu_{B}=\frac{1}{\left(\sigma_{\pi}^{-2}+n\right)}\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}{X_{i}}\right)$
Therefore, by Equation 10, supposing that $M=\beta$, and rearranging we have that the prior mean is required to be in the closed ball around the empirical mean
$\mu_{\pi}\in\overline{\mathrm{B}\left(\overline{X_{1}^{n}},\frac{2}{n}+\sigma_{\pi}^{2}\right)}$
in order for the Bayesian posterior to perform similarly to the GVI posterior in terms of robustness to prior misspecification. Equation 11 would make the computation tedious without improving clarity; for illustrative purposes this suffices. Here, we can interpret this as requiring the prior to have its mean within one variance of the empirical mean plus an error term depending on the number of observations.
5 Proofs
We now provide the proofs of the theorems presented in Section 3.
5.1 Characterising the GVI Posterior through the Empirical Loss Minimiser
We first present the following auxiliary lemmata that will be instrumental in the proof of Theorem 1.
Lemma 1.
Let Assumptions 1, 2 and 3 hold. Suppose $\exists C\in\mathbb{R}$ and that $\exists P,Q\in\mathcal{Q}$ such that
$\mathbb{E}_{P}[L(X_{1}^{n},\,\theta)]\leq C,\;\mathrm{and}\;\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]>C+\frac{M}{n\beta},$
then $\forall\Pi\in\mathcal{G}$ it holds that $T_{n}(P)<T_{n}(Q)$.
Proof of Lemma 1.
Recall that $T_{n}(\nu)=n\mathbb{E}_{\nu}[L(X_{1}^{n},\theta)]+\frac{1}{\beta}D(\nu:\Pi)$ for some $\nu\in\mathcal{Q}$, then for any $P,Q\in\mathcal{Q}$ satisfying the assumption above, we have that:
$T_{n}(P)=n\mathbb{E}_{P}[L(X_{1}^{n},\theta)]+\frac{1}{\beta}D(P:\cdot)\leq nC+\sup_{\Pi\in\mathcal{G}}\frac{1}{\beta}D(P:\Pi)\leq nC+\frac{M}{\beta}$
$T_{n}(Q)=n\mathbb{E}_{Q}[L(X_{1}^{n},\theta)]+\frac{1}{\beta}D(Q:\cdot)>n(C+\frac{M}{n\beta})+\inf_{\Pi\in\mathcal{G}}\frac{1}{\beta}D(Q:\Pi)\geq nC+\frac{M}{\beta}$
hence $T_{n}(P)<T_{n}(Q)$ for all priors $\Pi\in\mathcal{G}$. ∎
The next lemma is more general and does not require any assumptions on the divergence.
Lemma 2.
Make Assumptions 1, 2 and 5. Suppose $\mathcal{Q}$ is closed with respect to weak convergence. Pick any $C\in\mathbb{R}$, then the subspace $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq C\}$ is tight.
Proof.
Fix $\varepsilon>0$, and we show that $\exists K\subset\Theta$ compact, depending only on $\varepsilon$ and the loss, such that $\forall\nu\in\mathcal{A}$ it holds $\nu K>1-\varepsilon$.
We may decompose the functional, for a compact set $K\subset\Theta$ to be specified later, into
$J_{L_{n}}(\nu)=\int_{K}L_{n}\,d\nu+\int_{K^{c}}L_{n}\,d\nu\geq\nu K\inf_{\theta\in K}L_{n}(\theta)+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\,.$
As the loss is lower bounded and lower semi–continuous it achieves a minimum on any compact set, so $\inf_{\theta\in K}L_{n}(\theta)=c\in\mathbb{R}$ is achieved for some $\theta$ and exists, and hence by the lower bound, there exists some global minimum $c^{\star}:=\inf_{\theta\in\Theta}L_{n}(\theta)$. Now, without loss of generality we may assume $c^{\star}<0$ otherwise we may pick 0 as a lower bound instead of $c^{\star}$ below, so that
$\nu K\inf_{\theta\in K}L_{n}(\theta)+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\geq c^{\star}+\nu K^{c}\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)\,.$
Suppose the contrapositive that $\mathcal{A}$ is not tight, then we show that the hypothesis $J_{L_{n}}(\nu)\leq C$ also fails. In particular, this implies $\nu K^{c}\geq\varepsilon$, so
$J_{L_{n}}(\nu)\geq c^{\star}+\varepsilon\inf_{\theta\in\Theta\backslash K}L_{n}(\theta)$
and additionally, the set $c^{\star},\frac{C-c^{\star}}{\varepsilon}]\subset\mathbb{R}$ is compact. Therefore, by the coerciveness of the loss there exists $K^{\prime}\subset\Theta$ compact such that $L_{n}:\Theta\backslash K^{\prime}\rightarrow\mathbb{R}\backslash[c^{\star},\frac{C-c^{\star}}{\varepsilon}]$, in words there exists some compact set $K^{\prime}$ such that the loss defined on its complement maps $\theta$ exclusively outside of $c^{\star},\frac{C-c^{\star}}{\varepsilon}]$. Note that $K^{\prime}$ does not depend on $\nu$. We may now define $K=K^{\prime}$ to see that if $\mathcal{A}$ is not tight then $\exists\nu\in\mathcal{A}$ for which $J_{L_{n}}(\nu)>C$. As $\varepsilon$ was arbitrary, the claim follows. ∎
Using these lemmata, we are now able to prove Theorem 1.
Proof of Theorem 1.
By Assumption 4, $\exists P_{n}^{\star}\in\mathcal{Q}$ such that
$P_{n}^{\star}\in\operatorname*{arg\,inf}_{Q\in\mathcal{Q}}\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)].$
Now let $C=\mathbb{E}_{P_{n}^{\star}}[L(X_{1}^{n},\,\theta)]$. Then for all $Q\in\mathcal{Q}$ that satisfy
$\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]>C+M(n\beta)^{-1},$
by Lemma 1 it must hold that $T_{n}(P_{n}^{\star})<T_{n}(Q)$, and hence $Q(d\theta)$ is not a minimiser of the GVI objective. Therefore, the only measures $Q^{\prime}\in\mathcal{Q}$ that are potential minimisers of the GVI objective for any prior $\Pi\in\mathcal{G}$ are those that satisfy:
$T_{n}(Q^{\prime})\leq T_{n}(P_{n}^{\star}).$
And since $T_{n}(P_{n}^{\star})\leq nC+\beta^{-1}M$, we have
$T_{n}(Q^{\prime})=n\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]+\beta^{-1}D(Q^{\prime}:\Pi)\leq nC+\beta^{-1}M$
$\implies n(\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]-C)\leq\beta^{-1}(M-D(Q^{\prime}:\Pi))$
$\implies\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]-\mathbb{E}_{P_{n}^{\star}}[L(X_{1}^{n},\,\theta)]\leq M(n\beta)^{-1}$
Therefore, all probability measures $Q^{\prime}\in\mathcal{Q}$ that satisfy the above bound are potential minimisers of $T_{n}(\cdot)$. This is precisely the set $\mathcal{R}^{\star}_{n}$. Furthermore, by Assumption 4 we know that $\forall Q^{\prime}\in\mathcal{Q}$ we have $\mathbb{E}_{Q^{\prime}}[L(X_{1}^{n},\,\theta)]\geq C$, so the LHS is non–negative, and since $P_{n}^{\star}$ is in this set, it is non–empty.
For the moreover part, we additionally suppose that $\mathcal{Q}$ is weak closed and that the loss is lower semi–continuous and coercive. It immediately follows that the functional
$J_{L_{n}}(Q):=\mathbb{E}_{Q}[L(X_{1}^{n},\,\theta)]$
is lower semi–continuous, note that this follows by the functional being over probability measures and the loss being lower bounded and lower semi–continuous [11, 4]. We then know by Lemma 2 that the sublevel sets of the functional $J_{L_{n}}(\cdot)$ are tight. Hence, by the lower semi–continuity of the functional, $\mathcal{Q}$ being weak closed, and Prokhorov’s theorem [8], the set $\mathcal{A}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)\leq k\}$ is weak compact.
Therefore, we may apply this to $\mathcal{R}_{n}^{\star}$ by picking $k=M(n\beta)^{-1}+\inf_{\nu\in\mathcal{Q}}J_{L_{n}}(\nu)$, which exists and is finite by Assumption 4, establishing that $\mathcal{R}_{n}^{\star}$ is weak compact as desired. ∎
5.2 Existence and Uniqueness of a Generalised Variational Inference Posterior
We first prove existence of GVI minimisers and then show that if a convex combination is in $\mathcal{Q}$ then it will also be in $\mathcal{R}_{n}^{\star}$ from which uniqueness then immediately follows.
Proof of Theorem 2.
By Theorem 1 we have that $Q_{n}\in\mathcal{R}_{n}^{\star}$ which is weak compact. Then, as the divergence is lower semi–continuous, we have that the functional
$T_{n}(\nu):=nJ_{L_{n}}(\nu)+\frac{1}{\beta}D(\nu:\Pi)$
is lower semi–continuous [11]. Then, as the loss is lower bounded, and the divergence is non–negative, it follows that $T^{\star}:=\inf_{\nu\in\mathcal{Q}}T_{n}(\nu)>$-$\infty$, and we may construct an arbitrary sequence $\ell_{k}:=T_{n}(\nu_{k})$, where $(\nu_{k})_{k}\subset\mathcal{R}_{n}^{\star}$, which satisfies $\ell_{k}\rightarrow T^{\star}$ as $k\rightarrow\infty$.
As $\mathcal{R}_{n}^{\star}$ is weak compact, and the functional $\nu\mapsto T_{n}(\nu)$ is lower semi–continuous, it follows that the sequence $(\nu_{k})_{k}$ is tight. Therefore, by Prokhorov’s theorem, after passing to a subsequence (without relabelling) we find that $\nu_{k}\overset{w^{\star}}{\rightarrow}\nu_{\star}$ for some $\nu_{\star}\in\mathcal{R}_{n}^{\star}$ and for which
$\liminf_{k}\;T_{n}(\nu_{k})\geq T_{n}(\nu_{\star})\,.$
Further note that $\liminf_{k}\,T_{n}(\nu_{k})=T^{\star}$ and as $T^{\star}$ is a global minimum of the GVI objective, it follows that $T_{n}(\nu_{\star})=T^{\star}$, so $\nu_{\star}\in\mathcal{Q}$ exists which minimises Equation 3. ∎
Now, we provide the proof of Corollary 3 from which Corollary 4 is an immediate consequence, as the GVI functional under the assumptions of Corollary 4 is strictly convex, and hence a minimiser needs to be a unique extreme point of $\mathcal{R}_{n}^{\star}$.
Proof of Corollary 3.
The forward direction is immediate, given that $\mathcal{R}^{\star}_{n}$ is a subset of $\mathcal{Q}$.
For the reverse direction, assume there exists $Q,P\in\mathcal{R}^{\star}_{n}$, and assume that for some $a\in(0,1)$ it holds that $aQ+(1-a)P\in\mathcal{Q}$. Then we have that, as $J_{L_{n}}(\cdot)$ is a linear functional,
$J_{L_{n}}(aQ$ $+(1-a)P)-J_{L_{n}}(P_{n}^{\star})$
$=$ $aJ_{L_{n}}(Q)+(1-a)J_{L_{n}}(P)-J_{L_{n}}(P_{n}^{\star})$
$=$ $a\left[\underbrace{J_{L_{n}}(Q)-J_{L_{n}}(P_{n}^{\star})}_{\leq M(n\beta)^{-1}}\right]+(1-a)\left[\underbrace{J_{L_{n}}(P)-J_{L_{n}}(P_{n}^{\star})}_{\leq M(n\beta)^{-1}}\right]$
$\leq$ $M(n\beta)^{-1}$
Hence, $aQ+(1-a)P\in\mathcal{R}^{\star}_{n}$ and the reverse implication holds. ∎
5.3 Support of the Prior is not required
Proof of Proposition 5.
Suppose that $\exists Q\in\mathcal{Q}$ such that both:
  1. 1.
    $Q$ and $\Pi$ are mutually singular, and
  2. 2.
    $T_{n}(Q)\leq nC+\beta^{-1}M$, for $C$ as in the proof of Theorem 1,
hold. Then, as by Assumption 3, and the definition of a divergence $0\leq D(Q:\Pi)\leq M$. Either $D(Q:\Pi)=M$, in which case $Q$ minimises the loss functional $J_{L_{n}}(\cdot)$, or $D(Q:\Pi)=k<M$, in which case $Q$ is still an element of $\mathcal{R}^{\star}_{n}$. Moreover, we have that
$nJ_{L_{n}}(Q)+\beta^{-1}D(Q:\Pi)=nJ_{L_{n}}(Q)+\beta^{-1}k\leq nC+\beta^{-1}M$
$\implies n\Bigl[J_{L_{n}}(Q)-J_{L_{n}}(P_{n}^{\star})\Bigr]\leq\beta^{-1}(M-k)$
as desired.∎
5.4 Asymptotics and Frequentist Consistency
In order to prove this result we need to collect some auxiliary results that will establish the final result.
5.4.1 Charactering the loss minimiser
We begin with a general result for $\mathcal{Q}=\mathcal{P}(\Theta)$.
Lemma 3.
Suppose that $f:\Theta\rightarrow\mathbb{R}$ is a $\mathcal{T}$–measurable, lower–bounded, lower semi–continuous, and coercive function. Any minimiser of the functional $J_{f}(\cdot)$ must put full measure on any measurable set containing the minimisers of the function $f(\theta)$.
Proof.
Let $\varepsilon>0$ and denote $N_{\varepsilon}:=\{\theta\in\Theta:f(\theta)\leq\inf_{\vartheta\in\Theta}f(\vartheta)+\varepsilon\}$. For ease of notation, we denote $\inf_{\theta\in A}f(\theta)\equiv f(A)$. We then note that for $\nu\in\mathcal{P}(\Theta)$
$J_{f}(\nu):=\int_{\Theta}fd\nu=\int_{N_{\varepsilon}}fd\nu+\int_{N_{\varepsilon}^{c}}fd\nu\geq f(\Theta)\,.$
As the function is lower bounded, lower semi–continuous and coercive, it follows that there exists some minimiser away from the boundary of $\Theta$, specifically it achieves its minimum on some compact subset of $\Theta$ away from the boundary, so $N_{0}:=\{\theta\in\Theta:f(\theta)=f(\Theta)\}$ is well defined and non–empty. Note that this is measurable as we may define it as a countable intersection of measurable sets.
Then for any measure $\nu\in\mathcal{P}(\Theta)$ for which $\nu N_{0}=1$ the lower bound above is achieved. Hence, $\forall\varepsilon>0$ and $\forall\mu\in\mathcal{P}(\Theta)$ for which $\mu N_{\varepsilon}<1$ holds, it follows that $J_{f}(\nu)<J_{f}(\mu)$. We may see this as
$J_{f}(\mu)$ $=\int_{N_{\varepsilon}}fd\mu+\int_{N_{\varepsilon}^{c}}fd\mu$
$=\int_{\{f\leq f(\Theta)+\varepsilon\}}fd\mu+\int_{\{f>f(\Theta)+\varepsilon\}}fd\mu$
$\geq\mu N_{\varepsilon}\cdot f(\Theta)+\mu N_{\varepsilon}^{c}\cdot(f(\Theta)+\varepsilon)$
$=\mu N_{\varepsilon}f(\Theta)+(1-\mu N_{\varepsilon})(f(\Theta)+\varepsilon)$
$=f(\Theta)+\varepsilon(1-\mu N_{\varepsilon})>f(\Theta)$
establishing the claim. ∎
Therefore, we arrive at the natural conclusion that a measure minimising the expected loss, $P_{n}^{\star}$, is a Dirac delta measure at minimisers of the loss function.
Lemma 4.
Make Assumptions 1, 2 and 5, and let $\mathcal{Q}=\mathcal{P}(\Theta)$. Then $(P^{\star}_{n})_{n}$ will be a sequence of Dirac measures placing full measure on the set of points minimising the corresponding loss. As a consequence we have that $J_{L_{n}}(P_{n}^{\star})=L_{n}(\Theta)$.
Proof.
By Lemma 3 we have that the measure minimising $J_{L_{n}}(\cdot)$ will have full measure on the set $N^{n}_{0}:=\{\theta\in\Theta:L(X_{1}^{n},\,\theta)=L_{n}(\Theta)\}$. Since any such measure achieves the lower bound on the functional $J_{L_{n}}(\cdot)$ we may pick any $\theta_{n}^{\star}\in N^{n}_{0}$ and let $P^{\star}_{n}:=\delta_{\theta^{\star}_{n}}$. Therefore, $\delta_{\theta^{\star}_{n}}N_{0}^{n}=1$ as required. ∎
As a result we may characterise convergence of the expected loss minimising measures by the sets $(N_{0}^{n})_{n}$, however noting that the measures are not necessarily unique, even when the space we are minimising over is convex. We may now make the following observation.
Lemma 5.
Make Assumptions 1, 2, 5 and 7 and suppose $\mathcal{Q}=\mathcal{P}(\Theta)$. Let $A\in\mathcal{T}$ be any set for which $L(A)>L(\Theta)$. Suppose further that for any such $A$ it almost surely holds that
$\liminf_{n}\;L_{n}(A)>L(\Theta)\,$,
then $P_{0}$ almost surely $P_{n}^{\star}A\rightarrow 0$ as $n\rightarrow\infty$.
Proof.
For any such set $A$, there exists an $\varepsilon>0$ such that the sets $N_{\varepsilon}$ and $A$ are disjoint. In particular, by Lemma 3 $\forall n\in\mathbb{N}$ $P_{n}^{\star}N_{\varepsilon}^{n}=1$ $\forall\varepsilon>0$. Therefore, it remains to be shown that eventually almost surely $\exists\varepsilon>0$ such that $A\cap N_{\varepsilon}^{n}=\emptyset$ for all $n$ sufficiently large which will imply the claim.
To see this, by the assumption it is almost sure that
$\liminf_{n}L_{n}(A)>L(\Theta)$
which implies we may pick $\varepsilon=1/2(\liminf_{n}L_{n}(A)-L(\Theta))>0$. Further, as $L_{n}\rightarrow L$ as $n\rightarrow\infty$ pointwise, and $\limsup_{n}L_{n}(\Theta)\leq L(\Theta)$ ($P_{0}$–a.s.) by Assumption 7, so for all $n$ sufficiently large it holds almost surely that $L_{n}(A)-L_{n}(\Theta)>0$. And hence, eventually $P_{0}$–a.s. $A$ and $N_{\varepsilon}^{n}$, with $\varepsilon$ as above, will be disjoint. ∎
5.4.2 A Concentration Result of the GVI Posterior
It is now clear that $\forall\varepsilon>0$ the sequence of loss minimisers $(P_{n}^{\star})_{n}$ concentrates on a set $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}$. Therefore, it seems natural in view of Theorem 6 that the GVI posterior must also concentrate on this set. By Theorem 2 these posteriors exist for all $n$ so we may reason about the behaviour of $(Q_{n})_{n}$. We may explore this rigorously as follows.
Lemma 6.
Make Assumptions 1, 5, 3, 2, 4, 6 and 7, and suppose that $\mathcal{Q}=\mathcal{P}(\Theta)$. Recall the definition of the set $N_{\varepsilon}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon\}$. Suppose that $\forall\varepsilon>0$ the losses $P_{0}$–a.s. satisfy
$\liminf_{n}\,L_{n}(N_{\varepsilon}^{c})>L(\Theta)\,.$
Then $Q_{n}N_{\varepsilon}\rightarrow 1$ as $n\rightarrow\infty$ $P_{0}$–a.s. for any $\varepsilon>0$.
Proof.
We know that $P_{n}^{\star}N_{\varepsilon}\rightarrow 1$ as $n\rightarrow\infty$ $P_{0}$–a.s. $\forall\varepsilon>0$ by Lemma 5. By Theorem 1 we have that $\forall n$ the GVI posteriors $Q_{n}\in\mathcal{R}{{}_{n}^{\star}}:=\{\nu\in\mathcal{Q}:J_{L_{n}}(\nu)-J_{L_{n}}(P^{\star}_{n})\leq\frac{M}{\beta n}\}$.
Fix $\varepsilon>0$. Recall that by Lemma 4 we have $J_{L_{n}}(P_{n}^{\star})=L_{n}(\Theta)$, and by Assumption 7 $\limsup_{n}L_{n}(\Theta)\leq L(\Theta)$ (a.s.). By assumption we have that eventually $L_{n}(N_{\varepsilon}^{c})>L(\Theta)$ a.s. for all $n$ sufficiently large and therefore $\liminf_{n}L_{n}(N_{\varepsilon})=\liminf_{n}L_{n}(\Theta)$ $P_{0}$–a.s.
By Theorem 1 we may write for all $n\in\mathbb{N}$
$\frac{M}{\beta n}$ $\geq J_{L_{n}}(Q_{n})-J_{L_{n}}(P^{\star}_{n})$
$=\int_{N_{\varepsilon}}L_{n}dQ_{n}+\int_{N_{\varepsilon}^{c}}L_{n}dQ_{n}-L_{n}(\Theta)$
$\geq Q_{n}N_{\varepsilon}\cdot L_{n}(N_{\varepsilon})+Q_{n}N_{\varepsilon}^{c}\cdot L_{n}(N_{\varepsilon}^{c})-L_{n}(\Theta)$
$=Q_{n}N_{\varepsilon}\left(L_{n}(N_{\varepsilon})-L_{n}(\Theta)\right)+Q_{n}N_{\varepsilon}^{c}\left(L_{n}(N_{\varepsilon}^{c})-L_{n}(\Theta)\right)\geq 0\,.$ (12)
This is strictly non–negative $\forall n$ as $L_{n}(A)\geq L_{n}(\Theta)$ for all sets $A\in\mathcal{T}$. Then eventually $P_{0}$–a.s., $L_{n}(N_{\varepsilon}^{c})>L(\Theta)$, and it suffices to show the behaviour of $Q_{n}$} on $N_{\varepsilon}^{c}$. Therefore, $P_{0}$ almost surely
$L_{n}(N_{\varepsilon})-L_{n}(\Theta)\geq L_{n}(\Theta)-L_{n}(\Theta)=0\,,\qquad\liminf_{n}L_{n}(N_{\varepsilon}^{c})-L_{n}(\Theta)>0\,.$
As $M(\beta n)^{-1}$ converges to 0 uniformly, and we have the non–negativity of Equation 12, we conclude that $Q_{n}N_{\varepsilon}^{c}\rightarrow 0$ as $n\rightarrow\infty$ (a.s.), establishing the claim. ∎
We may compare this result with Miller [28] who draws the same conclusion for the Generalised Bayesian posteriors, i.e. GVI with $D=D_{KL}$ and $Q=\mathcal{P}(\Theta)$, with the caveat that there is the requirement on the prior, namely that $\Pi N_{\varepsilon}>0$ which we do not have here. Hence, we may concentrate on areas of the space $\Theta$ which do not have support under the prior, and on which the prior is badly behaved.
5.4.3 Establishing the Concentration of Generalised Variational Inference Posteriors
We may now prove Theorem 6 using these previous lemmas where we first make the following observation. When comparing Lemma 6 with Lemma 5, we note that $A=N_{\varepsilon}^{c}$ is such a set for which $L(A)>L(\Theta)$, and for all such sets $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$ there is some $\varepsilon>0$ small enough such that $A\subset N_{\varepsilon}^{c}$.
Proof of Theorem 6.
Fix $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$, then we may pick an $\varepsilon>0$ small enough such that $L(A)>L(\Theta)+\varepsilon$. In particular we may pick $\varepsilon=1/2(L(A)-L(\Theta))>0$, then $A\subset N_{\varepsilon}^{c}$.
Since $\mathcal{Q}\subset\mathcal{P}(\Theta)$ allows for Dirac measures we have that for all $n$ sufficiently large $\inf_{\nu\in\mathcal{Q}}J_{L_{n}}(\nu)=L_{n}(\Theta)$. By Lemma 2 and reasoning similar to the proof of Theorem 2 for each $n$ $\exists(\nu^{n}_{k})_{k}\subset\mathcal{Q}$ s.t. after passing to a subsequence in $k$ (without relabelling), $\nu_{k}^{n}\overset{w^{\star}}{\rightarrow}\nu_{\star}^{n}\in\mathcal{Q}$ as $k\rightarrow\infty$ for which $J_{L_{n}}(\nu_{\star}^{n})=L_{n}(\Theta)$ $P_{0}$–a.s. It follows Lemma 6 applies and hence $Q_{n}N_{\varepsilon}\rightarrow 1$ (a.s.) so $Q_{n}A\rightarrow 0$ (a.s.).
For the moreover part, we assume that $\exists\theta^{\star}\in\Theta$ uniquely minimising $L:\Theta\rightarrow\mathbb{R}$. Therefore, the set $N_{0}:=\{\theta\in\Theta:L(\theta)=L(\Theta)\}$ is a singleton namely $N_{0}=\{\theta^{\star}\}$ and hence $\delta_{\theta}^{\star}$ minimises the functional $J_{L}(\cdot)$. As per the previous $Q_{n}N_{\varepsilon}\rightarrow 1$ $\forall\varepsilon>0$ $P_{0}$–a.s. We may establish the following for sets $A\in\mathcal{T}$:
  1. 1.
    If $\theta^{\star}\in A$ and $L(A)=L(\Theta)$, then $\delta_{\theta^{\star}}A=1$ so $Q_{n}A\rightarrow 1$, and
  2. 2.
    If $\theta^{\star}\notin A$ and $L(A)>L(\Theta)$, then $\delta_{\theta^{\star}}A=0$ so $Q_{n}A\rightarrow 0$
[…]
$\left|\int_{\Theta}f\,dQ_{n}-\int_{\Theta}f\,d\delta_{\theta^{\star}}\right|\leq 2\varepsilon+\sum_{j}|\alpha_{j}|\,|Q_{n}(A_{j})-\delta_{\theta^{\star}}(A_{j})|$
We may pick $(A_{j})$ such that $\exists k$ for which $\delta_{\theta^{\star}}\in A_{k}$, and $\forall j\neq k$ $L(A_{j})>L(\Theta)$. Hence, the above applies and for all $n$ sufficiently large we have that the entire expression is eventually $P_{0}$–a.s. less than $3\varepsilon$. As $\varepsilon$ was arbitrary the result follows. ∎
Remark.
We have required that $\mathcal{Q}$ allows for Dirac–delta measures here, which we may satisfy by picking $\mathcal{Q}$ the set of Gaussian measures and define the Dirac Delta measures as a Gaussian with 0 covariance; see also Appendix A.
5.5 Rate of Convergence
Recall $N_{\varepsilon}:=\{\theta:L(\theta)\leq L(\Theta)+\varepsilon\}$ for which we know by Theorem 6 that $Q_{n}N_{\varepsilon}^{c}\rightarrow 0$. Now we instead pick a sequence $\varepsilon_{n}\rightarrow 0$ that is non–increasing, and positive, and we aim to show that if $\varepsilon_{n}$ decays sufficiently slowly, then we still have that $Q_{n}N_{\varepsilon_{n}}^{c}\rightarrow 0$. As $N_{\varepsilon_{n}}^{c}\rightarrow\Theta$, hence it will fail that $\liminf_{n}L_{n}(N_{\varepsilon_{n}})>L(\Theta)$. The preceding theory suggests that this should hold if $n\varepsilon_{n}\rightarrow\infty$. In fact, we may now show this under the assumptions of Theorem 7.
Proof of Theorem 7.
We may show this by demonstrating that $Q_{n}N_{\varepsilon_{n}}^{c}\rightarrow 0$. By the proof of Theorem 6 we may apply Equation 12 of Lemma 6, such that $\forall n\in\mathbb{N}$ sufficiently large
$\frac{M}{\beta}n^{-1}$ $\geq J_{L_{n}}(Q_{n})-J_{L_{n}}(P_{n}^{\star})$
$\geq Q_{n}N_{\varepsilon_{n}}(L_{n}(N_{\varepsilon_{n}})-L_{n}(\Theta))+Q_{n}N_{\varepsilon_{n}}^{c}(L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta))\geq 0\,.$
Further we have that
$N_{\varepsilon_{n}}:=\{\theta\in\Theta:L(\theta)\leq L(\Theta)+\varepsilon_{n}\}\rightarrow\{\theta\in\Theta:L(\theta)=L(\Theta)\}=:N_{0}$
which is well defined as the loss is lower semi–continuous. Then, by assumption on the loss and as it suffices to show the behaviour of $Q_{n}$ on $N_{\varepsilon_{n}}^{c}$, it follows that $L_{n}(N_{\varepsilon_{n}})\geq L_{n}(\Theta)$ and
$L_{n}(N_{\varepsilon_{n}})-L_{n}(\Theta)\geq L_{n}(\Theta)-L_{n}(\Theta)=0\,.$
So we may discard the term $Q_{n}N_{\varepsilon_{n}}(L_{n}(N_{\varepsilon_{n}})-L_{n}(\Theta))$ to get
$\frac{M}{\beta}n^{-1}\geq Q_{n}N_{\varepsilon_{n}}^{c}(L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta))\geq 0\,.$
We now need to show that $Q_{n}N_{\varepsilon_{n}}^{c}\rightarrow 0$. Taking these together, eventually $P_{0}$–a.s. $n$ is large enough for $\varepsilon_{n}$ to be sufficiently small such that we have $L_{n}(N_{\varepsilon_{n}}^{c})-L_{n}(\Theta)\gtrsim\varepsilon_{n}$. Therefore it follows that
$\frac{M}{\beta}n^{-1}$ $\gtrsim\varepsilon_{n}Q_{n}N_{\varepsilon_{n}}^{c}$
$\frac{M}{\beta}$ $\gtrsim n\varepsilon_{n}\cdot Q_{n}N_{\varepsilon_{n}}^{c}\,.$
Then, as $n\varepsilon_{n}\rightarrow\infty$, it must follow that $P_{0}$–a.s. $Q_{n}N_{\varepsilon_{n}}^{c}\rightarrow 0$ as otherwise, eventually
$n\varepsilon_{n}\gtrsim\frac{M}{\beta\cdot Q_{n}N_{\varepsilon_{n}}^{c}}$
$P_{0}$ almost surely, which would contradict $Q_{n}\in\mathcal{R}_{n}^{\star}$. ∎
5.6 Extensions
We collect here the proofs of the results from Section 3.6.
5.6.1 Existence and Uniqueness
Since Lemma 2 applies regardless of the divergence, with some more work we arrive at the conclusion of Proposition 8. In proving existence, we follow the ideas of [29] where they prove existence of minimisers for Kullback–Leibler divergence minimisation to some probability measure over a subset of possible probability measures over infinite dimensional Polish spaces.
Proof of Proposition 8.
As the loss is lower bounded, we may let this constant be $c:=\inf_{\theta}L(X_{1}^{n},\,\theta)$. Suppose without loss of generality that this constant is negative. Then we claim that $\forall C>c$ the sublevel sets of the GVI functional $T_{n}(\cdot)$ are weak compact. As the loss and the divergence are lower semi–continuous, it follows that $\nu\mapsto J_{L_{n}}(\nu)$ is lower semi–continuous and so $\nu\mapsto T_{n}(\nu):=J_{L_{n}}(\nu)+D(\nu:\Pi)$ is lower semi–continuous [11]. Hence, by Prokhorov’s theorem we need to show that sublevel sets are tight. This is immediate if $\Theta$ is compact, so we now show the result for $\Theta$ that are not.
Define $\mathcal{A}:=\{\nu\in\mathcal{Q}:T_{n}(\nu)\leq C\}$, then we aim to show that $\forall\varepsilon>0$ $\exists K\in\mathcal{T}$ compact, depending only on $\varepsilon$, such that $\forall\nu\in\mathcal{A}$ $\nu K>1-\varepsilon$.
By Lemma 2 we have that the sublevel sets of the functional $J_{L_{n}}(\cdot)$ are tight which is sufficient to show that $\mathcal{A}$ is tight.
[…]
Having shown that the sublevel sets of the GVI functional are weak compact, and as the functional is lower semi–continuous, we may now by the definition of the infimum, construct a sequence $(\ell_{k})_{k\in\mathbb{N}}\subset\mathbb{R}$ converging downwards to $\ell^{\star}:=\inf_{\nu\in\mathcal{Q}}T_{n}(\nu)$ defined via the corresponding elements $(\nu_{k})_{k}\subset\mathcal{A}$, i.e. $\ell_{k}:=T_{n}(\nu_{k})$. Since both the divergence and the loss are lower bounded, so is the GVI functional, so $\ell^{\star}>$-$\infty$. By the claim we find that $(\nu_{k})_{k}$ is a tight sequence of probability measures, so after passing to a subsequence (without relabelling) we must have that $\nu_{k}\overset{w^{\star}}{\rightarrow}\nu_{\star}$ for some $\nu_{\star}\in\mathcal{Q}$, and for which it holds that
$\liminf_{k}T_{n}(\nu_{k})\geq T_{n}(\nu_{\star})$
by the lower semi–continuity. As $T_{n}(\nu_{k})\rightarrow\ell^{\star}$, $\liminf_{k}T_{n}(\nu_{k})=\ell^{\star}$, and as $T_{n}(\nu_{\star})\geq\ell^{\star}$, we conclude that $T_{n}(\nu_{\star})=\ell^{\star}$ which is a global minimum of the GVI functional, so such a minimiser $\nu_{\star}\in\mathcal{Q}$ of Equation 3 exists.
For the uniqueness part, suppose that there exist two minimisers, $\nu_{1}$ and $\nu_{2}$. Then, as $\mathcal{Q}$ is convex we may pick $a\in(0,1)$ and let
$\nu:=a\nu_{1}+(1-a)\nu_{2}$ (13)
such that
$T_{n}(\nu)$ $=J_{L_{n}}\bigl(a\nu_{1}+(1-a)\nu_{2}\bigr)+D\bigl(a\nu_{1}+(1-a)\nu_{2}:\Pi\bigr)$
$<aJ_{L_{n}}\bigl(\nu_{1}\bigr)+(1-a)J_{L_{n}}\bigl(\nu_{2}\bigr)+aD(\nu_{1}:\Pi)+(1-a)D(\nu_{2}:\Pi)$
$=aT_{n}(\nu_{1})+(1-a)T_{n}(\nu_{2})$
where the inequality follows by the convexity of the divergence. But as $\nu_{1}$ and $\nu_{2}$ are both minimisers $T_{n}(\nu_{1})=T_{n}(\nu_{2})$ so $\nu$ would further lower the GVI objective, so these cannot be minimisers. Hence, we conclude that the only way Equation 13 may hold is if $\nu_{1}=\nu_{2}$ (in the sense of measures), and the minimiser may not be written as a convex combination of elements of $\mathcal{Q}$ so this minimiser is an extreme point, and unique. ∎
5.6.2 Relaxation of $M$ and $\beta$
Proof of Proposition 9.
The result is immediate from the corresponding proofs of Theorems 6 and 7. ∎
5.6.3 Concentration of GVI Posterior Measures under Unbounded Divergences
Proof of Theorem 10.
By Proposition 8, $Q_{n}$ exists under unbounded divergences. Then, for any such sequence $(\nu_{n})\subset\mathcal{Q}$ it follows by assumption that $P_{0}$ almost surely
$\frac{1}{n}T_{n}(\nu_{n})=J_{L_{n}}(\nu_{n})+(n\beta)^{-1}D(\nu_{n}:\Pi)\rightarrow L(\Theta)\,.$
Pick any such set $A\in\mathcal{T}$ for which $L(A)>L(\Theta)$. It follows that $P_{0}$–a.s.
$\frac{1}{n}T_{n}(\nu_{n})\geq\frac{1}{n}T_{n}(Q_{n})\geq J_{L_{n}}(Q_{n})=\int_{A}L_{n}(\theta)Q_{n}(d\theta)+\int_{A^{c}}L_{n}(\theta)Q_{n}(d\theta)\geq L_{n}(\Theta)\,.$
Note also that
$\int_{A}L_{n}(\theta)Q_{n}(d\theta)+\int_{A^{c}}L_{n}(\theta)Q_{n}(d\theta)\geq Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(A^{c})\geq L_{n}(\Theta)\,.$
By the lower semi–continuity of $L_{n}$ we have that $\exists\theta\in A^{c}$ for which $L(\theta)=L(\Theta)$ and therefore $L_{n}(A^{c})\rightarrow L(\Theta)$ $P_{0}$–a.s. Therefore,
$Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(A^{c})\geq Q_{n}A\>L_{n}(A)+Q_{n}A^{c}\,L_{n}(\Theta)\geq L_{n}(\Theta)$
Taking limits, it follows $P_{0}$ almost surely that $J_{L_{n}}(Q_{n})\rightarrow L(\Theta)$, and as $\liminf_{n}L_{n}(A)>L(\Theta)$ $P_{0}$–a.s. by assumption, it must follow that $Q_{n}A\rightarrow 0$ with $P_{0}$ probability 1. ∎
For unbounded divergences we will generally not have that $J_{L_{n}}(\nu_{n})=L_{n}(\Theta)$ at any finite $n$ as any atomic measure will have infinite divergence to any non–atomic prior. Therefore, we are unable to employ the same methodology in proving rates of convergence for GVI posterior measures under unbounded divergences as in the proof of Theorem 7.
6 Concluding Remarks
In this paper we have examined the asymptotic behaviour of Generalised Variational Inference measures on infinite dimensional hypothesis spaces with particular focus on prior misspecification. We have established frequentist consistency and rates of convergence to minimisers of $L(\theta)$ when using bounded divergences allowing for consistency regardless of the prior measure $\Pi\in\mathcal{G}$; see Figure 1. In fact, supposing that the divergence $D$ is everywhere bounded enables the choice $\mathcal{G}=\mathcal{P}(\Theta)$ hence including priors that are severely misspecified and result in inconsistent Bayesian posteriors.
It is worth noting that these results have further impact to Federated GVI [27], the study of GVI posteriors over a network of clients with private, local data sets. In particular such FedGVI posteriors will exist under the assumptions presented here through the different choices of loss functions and priors. Moreover, a global prior of some uniformed server, given a bounded divergence, may not skew the global posterior after incorporating local data towards unfavourable hypotheses that do not agree.
Finally, while we were able to derive existence and uniqueness of GVI posteriors under unbounded divergences, rates of convergence of GVI posterior measures under unbounded divergences remains an open challenge.
Appendix A Supplementary Definitions
As we expect this paper to be of interest also for researchers who are not as familiar with the intricacies of measure theory we recall some basic facts and definitions used throughout the paper without explicit statements.
We have used several modes of convergence of measures throughout the paper, we summarise these below.
Definition 2.
Let $(\mu_{n})_{n}$ and $\mu$ be Borel probability measures on the measurable space $(\Theta,\mathcal{T})$. We have the following notions of convergence.
  1. 1.
    If $\int_{\Theta}fd\mu_{n}\rightarrow\int_{\Theta}fd\mu$ $\forall f\in C_{b}(\Theta)$, then we say $\mu_{n}\overset{w^{\star}}{\rightarrow}\mu$, i.e. in the weak topology.
  2. 2.
    If $\mu_{n}A-\mu A\rightarrow 0$ $\forall A\in\mathcal{T}$, then $\mu_{n}\overset{sw}{\rightarrow}\mu$, i.e. setwise.
  3. 3.
    If $\mu_{n}-\mu||_{TV}\rightarrow 0$, then $\mu_{n}\overset{TV}{\rightarrow}\mu$, i.e. in total variation.
Here $C_{b}(\Theta)$ denotes the space of bounded, continuous functions mapping $\Theta$ to $\mathbb{R}$. Note, convergence in total variation implies setwise convergence, and setwise convergence in turn implies weak convergence, however the converse do not hold in general.
Definition 3.
A function $f:X\rightarrow Y$ between two topological spaces $(X,\mathcal{B})$ and $(Y,\mathcal{C})$ is called coercive if for every compact $K_{Y}\subset Y$ there exists a compact set $K_{X}\subset X$ such that the elements of $K_{X}^{c}$ map exclusively outside $K_{Y}$, that is
[…]
Definition 4.
We say that $\mathcal{Q}$ ‘allows for Dirac measures’ if (i) $\mathcal{Q}$ is closed with respect to weak convergence of probability measures, and (ii) if for all coercive, lower bounded and lower semi–continuous functions $f:\Theta\rightarrow\mathbb{R}$ it holds that
$\inf_{\nu\in\mathcal{Q}}\;J_{f}(\nu)=\inf_{\theta\in\Theta}\;f(\theta)\,.$
[…]
$m:=\int_{\mathcal{H}}\theta\nu(d\theta)\,,\quad\mathrm{and}\quad(\theta_{1},C\theta_{2})=\int_{\mathcal{H}}(\theta,\theta_{1})(\theta,\theta_{2})\nu(d\theta)\quad\forall\theta_{1},\theta_{2}\in\mathcal{H}$
where the first is a Bochner integral (see [29]).
Appendix B Classical Robustness to Prior Misspecification
We briefly discuss the relation of our work to robustness to prior misspecification in the sense of Bayesian sensitivity analysis [6, 17]. This line of work has examined the stability of Bayesian posteriors under small perturbations of the prior with respect to some other measure, for instance through additive Huber contamination as mentioned in Section 2.2. In particular, this has been used to select a robust prior for Bayes which leads to optimal decision making under potential perturbations of the prior measure. In this paper however, we aim to answer the reverse: under what conditions can GVI lead to optimal decision making under ill chosen priors.
Local sensitivity of posteriors generated with $\varepsilon$–contaminated priors, for instance by the additive Huber contamination model $\mathcal{G}=\{\Pi_{\varepsilon}:=(1-\varepsilon)\Pi_{0}+\varepsilon G:\varepsilon\in(0,1)\}$, is often studied [17] through the Gâteux differential
$s(\Pi_{0},G;X_{1}^{n})=\lim_{\varepsilon\downarrow 0}\cfrac{d(Q^{n}_{0}:Q^{n}_{\varepsilon})}{d(\Pi_{0}:\Pi_{\varepsilon})}$
where $d(\cdot:\cdot)$ is some distance or divergence measure; Gustafson and Wasserman [17] for instance use the total variation distance. Here $\Pi_{0}$ is the uncontaminated prior, $\Pi_{\varepsilon}$ the $\varepsilon$–contaminated prior in the direction of $G$, and $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ the respective Bayes posteriors. The unrestricted choice of a divergence other than the Kullback–Leibler in GVI generally prohibits explicit expressions of $Q^{n}_{0}$ and $Q^{n}_{\varepsilon}$ that demonstrate the the exact dependence on the prior, which makes this approach impracticable for our purposes.
This presents an interesting theoretical question to the stability of Bayesian posteriors with respect to prior misspecification and suggests fascinating future directions for understanding uncertainty about the hypotheses in GVI given finite observations. The work in [6, 17] is concerned with small perturbations in the prior within a restricted class of measures $\mathcal{G}$. Often, restrictions are placed on this set in terms of how much the prior may vary; in [12] the priors considered are those close to the original prior $\Pi_{0}$ as $\varepsilon$ tends to zero. The present paper however is concerned with severe prior misspecification, even cases where we allow $\mathcal{G}=\mathcal{P}(\Theta)$, as well as the asymptotic dynamics of GVI posterior measures under such misspecification.
Appendix C Non–Robustness in Bayesian Inference
Continuing the illustrative example from Sections 2.3 and 4.2, with $X_{1}^{n}\overset{iid}{\sim}\mathcal{N}(\theta,\sigma_{p}^{2})$, we get the GVI objective. Using $\beta=1$ and loss $L(X_{1}^{n},\,\theta)=-\frac{1}{n}\log p_{\theta}(X_{1}^{n})$, $\mathbb{E}_{q(\theta)}[nL(X_{1}^{n},\,\theta)]$ is
$\mathbb{E}_{q(\theta)\sim\mathcal{N}(\mu_{q},\sigma_{q}^{2})}[-\log p_{\theta}(X_{1}^{n})]=\mathbb{E}_{q(\theta)}\left[-\log{(2\pi\sigma_{p}^{2})^{-\frac{n}{2}}}\exp\left\{-{\sum_{i=1}^{n}}\frac{(\theta-X_{i})^{2}}{2\sigma_{p}^{2}}\right\}\right]$
$\frac{n}{2}\log 2\pi\sigma_{p}^{2}+\frac{1}{2\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})^{2}+\frac{n\sigma_{q}^{2}}{2\sigma_{p}^{2}}$
Therefore, we have a closed form for $R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ as
$\frac{n}{2}\log 2\pi\sigma_{p}^{2}+\frac{1}{2\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})^{2}+\frac{n\sigma_{q}^{2}}{2\sigma_{p}^{2}}+\log\frac{\sigma_{\pi}}{\sigma_{q}}+\frac{(\mu_{q}-\mu_{\pi})^{2}+\sigma_{q}^{2}}{2\sigma_{\pi}^{2}}-\frac{1}{2}$
which is convex in both $\mu_{q}$ and $\sigma_{q}$. The minimisers of this objective are hence
$\nabla_{\mu_{q}}R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ $=\frac{1}{\sigma_{p}^{2}}\sum_{i=1}^{n}(\mu_{q}-X_{i})+\frac{\mu_{q}-\mu_{\pi}}{\sigma_{\pi}^{2}}=0$
$\nabla_{\sigma_{q}}R(D_{KL},-\frac{1}{n}\log p_{\theta},q)$ $=\frac{n\sigma_{q}}{\sigma_{p}^{2}}-\frac{1}{\sigma_{q}}+\frac{\sigma_{q}}{\sigma_{\pi}^{2}}=0$
Solving for $\sigma_{q}$ and $\mu_{q}$ gives the well known solution
$\sigma_{q}^{2}=\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)^{-1},\,$ $\mu_{q}=\cfrac{\left(\frac{\mu_{\pi}}{\sigma_{\pi}^{2}}+\sum_{i=1}^{n}\frac{X_{i}}{\sigma_{p}^{2}}\right)}{\left(\sigma_{\pi}^{-2}+n\sigma_{p}^{-2}\right)}$
which shows that, for a finite $\sigma_{\pi}^{2}$, $\mu_{\pi}$ by itself can cause arbitrary misspecification.
{acks}
[Acknowledgments] The first author would like to direct correspondence to either Terje.Mildner@kellogg.ox.ac.uk or Terje.Mildner@gmail.com. {funding} TM and TD were supported by a UKRI Turing AI Acceleration Fellowship [EP/V02678X/1] and a Turing Impact Award from The Alan Turing Institute.